OpenStudy (erinweeks):

A basket contains 15 pieces of fruit: 7 apples, 5 oranges, and 3 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple?

OpenStudy (erinweeks):

A. 6/7 B. 7/45 C. 1/6 D. 4/5

OpenStudy (lgbasallote):

probablility jonas will get orange is 5/15 <--because 5 oranges out of 15 fruits probability Beth will get apple = 7/14 <-- 7 apples out of 14 fruits (because 1 fruit has already been taken) so the overall probability is \[\frac 5 {15} \times \frac 7{14}\] does that help?

OpenStudy (erinweeks):

but im not getting a right answer !

OpenStudy (lgbasallote):

simplify then

OpenStudy (lgbasallote):

\[\frac 5{15} = \frac 13\] \[\frac 7{14} = \frac 12\] do you get the answer now?

OpenStudy (erinweeks):

no i got 5/6 thats not a answer !

OpenStudy (lgbasallote):

how did you get 5/6?

OpenStudy (lgbasallote):

what is \[\frac 13 \times \frac 12\] ??

OpenStudy (erinweeks):

i dont know :/

OpenStudy (lgbasallote):

\[\frac 13 \times \frac 12 \implies \frac{1 \times 1}{3 \times 2}\]

OpenStudy (lgbasallote):

have you forgotten those lessons?

OpenStudy (erinweeks):

oh see i was adding !

OpenStudy (lgbasallote):

you usually multiply in probabilities

OpenStudy (erinweeks):

okay thanks !

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

1/6